Current operating income for Bay Area Cycles Co. is $26,000. Selling price per unit is $100, the contribution margin ratio is 25%, and fixed expense is $104,000. Required: Calculate Bay Area Cycle’s per unit variable expense and contribution margin. How many units are currently being sold? How many additional unit sales would be necessary to achieve operating income of $65,000?
Answer #1

a) Variable cost per unit = 100*75% = 75 per unit

Sales = (104000+26000)/.25 = $520000

Contribution margin = 520000*.25 = $130000

Current unit sold = 520000/100 = 5200 Units

b) Required unit sales = (104000+65000)/25 = 6760 Units

Additional unit sale to achieve operating income of $65000 = 6760-5200 = 1560 Units